gelphielvr:

(algebra 1) how do I figure out the y-intercept using this table? I already have the slope which is 5 but I'm not really sure how to get the y-intercept Table posted in the replies

gelphielvr:

1 attachment
SunFlower01:

do you know what desmos is ?

gelphielvr:

@sunflower01 wrote:
do you know what desmos is ?
heard of it but I have no clue what it is.

SunFlower01:

@sunflower01 wrote:
do you know what desmos is ?
well if u dont it helps with all dif math equations mostly ones that can be graphed

SunFlower01:

@gelphielvr wrote:
@sunflower01 wrote:
do you know what desmos is ?
heard of it but I have no clue what it is.
look it up and use the table by clicking "+" on the upper left corner

ThanosQC:

To solve this equation, first use slope formula (\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]) to find the slope. \[\frac{ 65-25 }{ 10-2 } = \frac{ 40 }{ 8 } = 5 = m\]. Next, use point slope form (\[y-y _{1}=m(x-x _{1})\] to find "b". \[y _{1}\] and \[x _{1}\] can be any coordinate on the table; y-25 = 5(x-2). Next, simply isolate and you have your equation in point-slope form: y-25 = 5(x-2) y-25 + 25 = 5x - 10 + 25 y = 5x + 15 b = 15, so the y-intercept of this table would be 15.